How do I put this problem into a calculator?
Problem: 0.71=log(x)
























Answer by Alan3354(69443)

How do I put this problem into a calculator?
Problem: 0.71=log(x)
-------------
IDK what kind of calculator you have.
============
0.71=log(x) is not a problem, it's an equation.
----
If you want to solve for x:
0.71=log(x)
10^0.71 = 10^(log(x)) = x
x =~ 5.1286